For the following set of cash flows,

  

Year Cash Flow
0 –$8,400           
1 3,600           
2 4,900           
3 6,200           

  

a. What is the NPV at a discount rate of 0 percent?

  

b. What is the NPV at a discount rate of 8 percent?

  

c. What is the NPV at a discount rate of 18 percent?
d. What is the NPV at a discount rate of 23 percent?
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=3,600+4,900+6,200=14700

NPV=Present value of inflows-Present value of outflows

=14700-8400

=$6300

b.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=3,600/1.08+4,900/1.08^2+6,200/1.08^3

=12456.05

NPV=Present value of inflows-Present value of outflows

=12456.05-8400

=$4056.05(Approx).

c.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=3,600/1.18+4,900/1.18^2+6,200/1.18^3

=10343.46

NPV=Present value of inflows-Present value of outflows

=10343.46-8400

=$1943.46(Approx).

d.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=3,600/1.23+4,900/1.23^2+6,200/1.23^3

=9497.42

NPV=Present value of inflows-Present value of outflows

=9497.42-8400

=$1097.42(Approx).